From GlueXWiki
Jump to: navigation, search

July 3, 2014 FDC+CDC meeting


  1. Commissioning plans
  2. Gas system
  3. DAQ (Beni)
  4. CDC status (Beni)
  5. FDC status fADC pulse window (Lubomir)
  6. Electronics (Chris, Nick)
  7. Engineering (Bill)
  8. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Naomi (CMU), Beni, Dave, Chris, Mike, Simon, Fernando, Eugene, Glenn, and Lubomir (JLab).

Readiness review

- Drafts (very preliminary) for the CDC and FDC commissioning were prepared for the readiness review (page linked above). For the FDC the plan so far is to use two gas mixtures, 90/10 and 40/60, the first one is better for the alignment. The target position(s) and turning on/off the magnetic field needs to be coordinated.

Gas system

- Last Friday, at ~8am, the power in the gas room was turned off for ~15 min due to a planned electrical work there. After turning it ON, the setting in the gas control panel were wrong (40mA, while the range is 4-20mA). As a result: (1) the flow was doubled and (2) the PLS didn't turned OFF the flow controllers when the mixing tanks reached 12psi. Soon the pressure reached 15psi and the two relief valves were continuously engaged for some time.

- After the meeting we installed an UPS (in preparation for the holiday weekend) there. Dave will contact the company to find a solution, at least to change the default settings to what we need.


- Beni: can read now all the 18 crates (CDC+FDC). Sometimes DAQ crashes fatally, so that you have to reboot the computer (suspect InfiniBand). Also somehow the bios on gluon51 failed; will be replaced.

- Beni made independent initialization for the FDC and CDC, now they can have different pulse windows.

CDC noise

- Fernando suspects the noise may come from the CDC itself, since it depends on the HV, and suggested increasing the CO2 content while keeping the same HV; the idea is to exclude the HV modules as a source of the noise. Eugene/Naomi: same effect will be if decrease/increase the HV by few hundred volts. Also it's difficult to explain with gas discharges the correlation we see in the noise (see plots from the last meeting). In any case, first Mike will study systematically where on the CDC we see the noise.

- Lubomir looked with a scope on the FDC and found noise with very similar pattern on several special channels. That was on the most upstream cathodes in the first and second package, but not in the other two, and always in cards that have both long and short strips (more susceptible to noise). It may indicate that the noise comes from the CDC or the mesh but one has to study this more systematically.

FDC update

- Lubomir started analyzing the fADC data from the last DAQ runs. The pedestal widths for the CDC are about a factor of 2 bigger than for the FDC: 20, 10 fADC units respectively. The widths in the FDC depend on the size of the strips.

- Important is that FDC sees many events with the CDC trigger; the explanation: most of the cosmic events are part of showers. The timing of the first maximum resembles the standard FDC drift time spectrum (histogram linked above). Beni will use this histogram to adjust the FDC pulse window from 20th to 100th sample, thus cutting the size of the files by 2. Before it was 0 to 160, as in the CDC.


- Fernando: two CAEN modules back from repair. Chris/Nick will test them before using in the Hall.


- Eugene: on the readiness review there was a remark that people stopped using plastic tubes for Fluorinert, only metal. Tim answered that the tube material was chosen to work for Fluorinert.

- Dave: put new mark on the Fluorunert level. Moved all the hardware needed for the cathode scan, the motor, controller and computer, in the ESB.


- Luke Robison, a graduate student from Northwestern University will work on FDC projects for ~2 months: noise studies, cathode scan.